### v1

@@ 3.2 Parametric regularity for uniform random fields
$\quad+e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}.$
@@
$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)(\cdot,t)\|_{H^{1}(\Omega)^{*}}$
$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w(\cdot,t)\|_{H^{1}(\Omega)}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w(\cdot,t)\|_{H^{1}(\Omega)}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w(\cdot,t)\|_{H^{1}(\Omega)}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w(\cdot,t)\|_{H^{1}(\Omega)},$ (3.15)
where we also used $\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w(\cdot,t)\|_{L^{2}(\Omega)}\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{C(I;L^{2}(\Omega))}\leq\theta_{2}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}$. This step to remove the dependence of one factor on $t$ is crucial before we integrate over $t\in I$.
Next we square both sides of (3.2), integrate over $t\in I$, and then take the square root. We can view the entire right-hand side of (3.2) as a sum of the form $\sum_{k}\alpha_{k}\,g_{k}(t)$, and so we use the estimate
$\bigg(\int_{I}\bigg(\sum_{k}\alpha_{k}\,g_{k}(t)\bigg)^{2}\,{\mathrm{d}}t\bigg)^{1/2}$ $=\bigg(\sum_{k}\sum_{k^{\prime}}\alpha_{k}\,\alpha_{k^{\prime}}\int_{I}g_{k}(t)\,g_{k^{\prime}}(t)\,{\mathrm{d}}t\bigg)^{1/2}$
$\leq\bigg(\sum_{k}\sum_{k^{\prime}}\alpha_{k}\,\alpha_{k^{\prime}}\bigg(\int_{I}g_{k}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}\bigg(\int_{I}g_{k^{\prime}}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}\bigg)^{1/2}$
$=\sum_{k}\alpha_{k}\,\bigg(\int_{I}g_{k}^{2}(t)\,{\mathrm{d}}t\bigg)^{1/2}.$
This leads to
$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{L^{2}(I;H^{1}(\Omega))}$
$\leq(a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}},$ (3.16)
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $=\sqrt{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}$
$\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(1+a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}}$
$\leq(1+a_{\max}+\kappa_{\max}+\lambda+f_{\max})\,\frac{\Theta_{\boldsymbol{\nu}}}{\min(a_{\min},\lambda-\kappa_{\max})}+\Theta_{\boldsymbol{\nu}},$
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $\leq\Phi\,\bigg[2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}w\|_{\mathcal{X}}$
$\qquad\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}w\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}w\|_{\mathcal{X}}\bigg],$ (3.17)
@@
$\Phi:=\frac{1+a_{\max}+\kappa_{\max}+\lambda+f_{\max}}{\min(a_{\min},\lambda-\kappa_{\max})}+1.$ (3.18)
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\qquad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}},$ (3.19)
@@
Our base step is $|{\boldsymbol{\nu}}|=1$ (rather than ${\boldsymbol{\nu}}={\boldsymbol{0}}$, as in typical regularity proofs) because our recursion (3.2) involves a sum over ${\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}$ which is empty if $|{\boldsymbol{\nu}}|=1$. Recall from Theorem 2.2 that we have the a priori bound $\|w\|_{\mathcal{X}}\leq C\,\|w_{0}\|_{L^{2}(\Omega)}$, where $\|w_{0}\|_{L^{2}(\Omega)}\leq 1$ since $0\leq w_{0}({\boldsymbol{x}})\leq 1$. For ${\boldsymbol{\nu}}={\boldsymbol{e}}_{k}$ for some $k\geq 1$, the recursion (3.2) gives
$\|\partial_{\boldsymbol{y}}^{{\boldsymbol{e}}_{k}}w\|_{\mathcal{X}}$ $\leq\Phi\,\big(2\,\beta_{k}\,\|w\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\beta_{k}\,\|w\|_{\mathcal{X}}^{2}\big)$
$\leq\Phi\,\big(2\,\beta_{k}\,C\,\|w_{0}\|_{L^{2}(\Omega)}+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\,\beta_{k}\,C^{2}\,\|w_{0}\|_{L^{2}(\Omega)}^{2}\big)\leq\frac{D\,\beta_{k}}{2},$
@@
$D:=2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big)\,\|w_{0}\|_{L^{2}(\Omega)}\leq 2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big)=:E.$ (3.20)
@@
$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$

### v2

@@ 3.2 Parametric regularity for uniform random fields
$\quad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}.$
The last factor in every term on the right-hand side can be upper bounded by $\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}$. Canceling out this common factor from both sides and grouping the first two terms on the right-hand side, we arrive at
$\frac{\min(a_{\min},\kappa_{\max})}{2\,e^{2\,\kappa_{\max}\,T}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}$
$\leq 2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}$
$\qquad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\,=:\,\Theta_{{\boldsymbol{\nu}}}.$ (3.16)
Time derivative. We return to (3.6) but remove the integration over $t$ and move all terms except the duality pair to the right-hand side. Then we apply the triangle, Cauchy–Schwarz and Hölder inequalities, together with $|2\,\kappa\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,u\,v+(f-\kappa)\,(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\,v|\leq(\kappa+f)\,|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u|\,|v|$ since $0\leq u\leq 1$. We obtain for a.e. $t\in I$,
$\big|\big\langle\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u),v\big\rangle_{H^{1}(\Omega)^{*},H^{1}(\Omega)}\big|$
$\leq a_{\max}\,\|\nabla(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+(\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\nabla(\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u)(\cdot,t)\|_{L^{2}(\Omega)}\,\|\nabla v\|_{L^{2}(\Omega)}+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|v\|_{L^{2}(\Omega)}$
$\quad+\kappa_{\max}\sum_{{\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}$
$\quad+\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u(\cdot,t)\|_{L^{2}(\Omega)}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u(\cdot,t)\|_{L^{4}(\Omega)}\,\|v\|_{L^{4}(\Omega)}.$
@@
$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)(\cdot,t)\|_{H^{1}(\Omega)^{*}}$ $\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u(\cdot,t)\|_{H^{1}(\Omega)}+2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u(\cdot,t)\|_{H^{1}(\Omega)}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u(\cdot,t)\|_{H^{1}(\Omega)}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u(\cdot,t)\|_{H^{1}(\Omega)},$ (3.17)
where we also used $\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u(\cdot,t)\|_{L^{2}(\Omega)}\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{C(I;L^{2}(\Omega))}\leq\theta_{2}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}$. This step to remove the dependence of one factor on $t$ is crucial before we integrate over $t\in I$.
Next we take the $L^{2}(I)$ norm of (3.2) and apply the triangle inequality to obtain
$\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$ $\leq(a_{\max}+\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}},$ (3.18)
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$ $=\sqrt{\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}^{2}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;H^{1}(\Omega)^{*})}^{2}}$
$\leq\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}+\|\tfrac{\partial}{\partial t}(\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u)\|_{L^{2}(I;H^{1}(\Omega)^{*})}$
$\leq(1+a_{\max}+\kappa_{\max}+f_{\max})\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{L^{2}(I;H^{1}(\Omega))}+\Theta_{\boldsymbol{\nu}}$
$\leq(1+a_{\max}+\kappa_{\max}+f_{\max})\,\frac{2\,e^{2\kappa_{\max}\,T}\,\Theta_{\boldsymbol{\nu}}}{\min(a_{\min},\kappa_{\max})}+\Theta_{\boldsymbol{\nu}},$
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$ $\leq\Phi\,\bigg[2\sum_{j\geq 1}\nu_{j}\,\beta_{j}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}u\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}u\|_{\mathcal{X}}$
$\qquad\quad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{m}}u\|_{\mathcal{X}}\,\|\partial_{\boldsymbol{y}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}u\|_{\mathcal{X}}\bigg],$ (3.19)
@@
$\Phi:=\frac{2\,e^{2\,\kappa_{\max}\,T}(1+a_{\max}+\kappa_{\max}+f_{\max})}{\min(a_{\min},\kappa_{\max})}+1.$ (3.20)
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\qquad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}},$ (3.21)
@@
Our base step is $|{\boldsymbol{\nu}}|=1$ (rather than ${\boldsymbol{\nu}}={\boldsymbol{0}}$, as in typical proofs of parametric regularity) because our recursion (3.2) involves a sum over ${\boldsymbol{0}}<{\boldsymbol{m}}<{\boldsymbol{\nu}}$ which is empty if $|{\boldsymbol{\nu}}|=1$. Recall from Theorem 2.2 that we have the a priori bound $\|u\|_{\mathcal{X}}\leq C\,\|u_{0}\|_{L^{2}(\Omega)}$, where $\|u_{0}\|_{L^{2}(\Omega)}^{2}\leq|\Omega|$ since $0\leq u_{0}({\boldsymbol{x}})\leq 1$. For ${\boldsymbol{\nu}}={\boldsymbol{e}}_{k}$ for some $k\geq 1$, the recursion (3.2) gives
$\|\partial_{\boldsymbol{y}}^{{\boldsymbol{e}}_{k}}u\|_{\mathcal{X}}$ $\leq\Phi\,\big(2\,\beta_{k}\,\|u\|_{\mathcal{X}}+\theta_{1}^{2}\,\theta_{2}\,\beta_{k}\,\|u\|_{\mathcal{X}}^{2}\big)$
$\leq\Phi\,\big(2\,\beta_{k}\,C\,\|u_{0}\|_{L^{2}(\Omega)}+\theta_{1}^{2}\,\theta_{2}\,\beta_{k}\,C^{2}\,\|u_{0}\|_{L^{2}(\Omega)}\,|\Omega|^{1/2}\big)\leq\frac{D\,\beta_{k}}{2},$
@@
$D:=E\,\|u_{0}\|_{L^{2}(\Omega)},\quad\mbox{with}\quad E:=2\,\Phi\,C\,\big(2+C\,\theta_{1}^{2}\,\theta_{2}\,|\Omega|^{1/2}\big).$ (3.22)
@@
$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$